# Uncertainty Management in the Construction of Knowledge Graphs: a Survey

### v1

@@ 6.1 Uncertain KG embedding models
Model Scoring function Evaluation task Dataset Year IIKE [41] TransE Link prediction NELL 2016 Triple classification FB15K (synthetic) URGE[62] Node Clustering PPI Matrix Factorization (proximity preservation) Node Classification DBLP 2017 k-NN search CKRL [151] KG Noise Detection TransE KG Completion FB15K (synthetic) 2018 Triple Classification GTransE [74] TransE KG Completion FB15K-237 (synthetic) 2019 NELL (synthetic) UKGE [20] Confidence prediction CN15K DistMult Relation fact ranking NL27K 2019 Relation fact classification PPI5K UOKGE [11] TransE Confidence Prediction CN15K 2019 Triples Classification SUKE [141] Link prediction CN15K DistMult Fact classification NL27K 2021 PPI5K BEURRE [21] Gumbel boxes Confidence prediction CN15K 2021 Relation fact ranking NELL27k PASSLEAF [22] Tail Entity Prediction PPI5K RotatE Confidence Prediction NL27K ComplEx CN15K 2021 DistMult WN18RR FB15K237 GMUC [161] Minimum Similarity Link Prediction NL27K 2021 Confidence Prediction ConfE [165] RESCAL Entity Type Noise Detection FB15kET (synthetic) 2021 Entity Type Prediction YAGO43k (synthetic) FocusE [111] Modifiable Scoring Layer Link Prediction (High-Valued Links) O*NET20K 2021 CN15K NL27K PPI5K WaExt [77] TransE CN15K 2022 TransH Link Prediction NL27K DistMult Triple Classification PPI5K ComplEx Wang et al. [142] Similarity Confidence Prediction CN15K 2022 Tail Entity Prediction NL27K MUKGE [91] Circular correlation Confidence Prediction CN15K 2024 Relation Fact Ranking NL27K Relation Fact Classification PPI5K
@@ 8.2 Fusion Approaches
Fusion methods are identified in Table 3, which provides an overview and indicates if the methods can handle some characteristics of data or data sources such as numerical data, data granularity, and dependency between data sources.
Model Task Modeling Awareness Datasets Year TI SQ Numerical Data Granularity Source dependence GTM [163] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ Wikipedia edit history of city population 2012 People biographies LTM [164] $\bullet$ $\bullet$ Probabilistic $\circ$ $\circ$ $\circ$ Book author 2012 Movie director Synthetic LCA [113] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ Books 2013 Population Stocks FantasySCOTUS KBT [34] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ Triples collected by Knowledge Vault 2015 Synthetic SLiMFAST [117] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ Stocks 2017 Demonstrations Crowds Genomics LFC [116] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ Digital mammography 2010 Breast MRI DOCS[166] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ ItemCompare 2016 4-Domain Yahoo QA SFV MDC [88] $\bullet$ $\bullet$ Representation Learning $\circ$ $\circ$ $\circ$ Dataset created from baobaozhido (crowdsourcing platform) 2017 POPACCU [33] $\circ$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ Books (from AbeBooks.com) 2012 Flight ACCU [32] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ Synthetic 2009 CRH [84] $\bullet$ $\bullet$ Loss Optimization $\bullet$ $\circ$ $\circ$ Weather Forecast (from Wunderground, HAM weather, World Weather Online) 2014 Stocks Flight Record Fusion [55] $\bullet$ $\circ$ Softmax Classifier $\bullet$ $\circ$ $\circ$ Flight 2020 Stock 1 Stock 2 Weather Address TruthFinder [159] $\bullet$ $\bullet$ Probabilistic $\circ$ $\bullet$ $\circ$ Book authors 2007 ASUMS [5] $\bullet$ $\bullet$ Belief functions $\circ$ $\bullet$ $\circ$ Synthetic 2016 People biographies TDH [72] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ BirthPlaces 2019 Heritages TKGC [63] $\bullet$ $\bullet$ Representation Learning $\bullet$ $\bullet$ $\circ$ Dataset built from [14] 2022 Probabilistic Subgraph of Freebase as prior knowledge (KG) OKELE [14] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ Synthetic 2020 FaitCrowd [92] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ SFV 2015 Dataset from crowdsourcing platform HYBRID [82] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ Book 2017 Synthetic CATD [83] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ $\circ$ City Population 2014 Biography Indoor Floorplan KDEm [140] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ Synthetic 2016 Population
@@
CRH [84] (Conflict Resolution on Heterogeneous Data) estimates the reliability of a source, it uses all types of data simultaneously instead of focusing on a single one, but requires availability of all data for each entity. To initiate the source reliability score, it first applies a simple conflict resolution method, such as majority or average voting.
@@
Then, they use the logarithm to facilitate the computation and to obtain the final scores. Although simple, these definitions integrate the influence of facts and take into account the dependence between data sources.
@@
OKELE [14] models the probability of a fact being true by a latent random variable following a Beta$(\beta_{0},\beta_{1})$ distribution, with $\beta_{1}$ corresponding to prior true count of the fact and $\beta_{0}$ corresponding to its prior false count. The quality of a data source is represented by its error variance $\omega_{s}$ that follows a scaled inverse chi-squared distribution Scale-inv-$\chi^{2}(v_{s},\tau_{s}^{2})$ representing the number of facts provided by the source $v_{s}$ with variance $\tau_{s}^{2}$. The authors argue that this distribution handles the effect of dataset size in the case of long-tail entities. The truth inference is performed by leveraging prior knowledge from existing KGs.

### v2

@@ 6.1 Uncertain KG embedding models
Model Scoring function Evaluation task Dataset Year IIKE [41] TransE Link prediction NELL 2016 Triple classification FB15K (synthetic) URGE[61] Node Clustering PPI Matrix Factorization (proximity preservation) Node Classification DBLP 2017 k-NN search CKRL [149] KG Noise Detection TransE KG Completion FB15K (synthetic) 2018 Triple Classification GTransE [73] TransE KG Completion FB15K-237 (synthetic) 2019 NELL (synthetic) UKGE [20] Confidence prediction CN15K DistMult Relation fact ranking NL27K 2019 Relation fact classification PPI5K UOKGE [11] TransE Confidence Prediction CN15K 2019 Triples Classification SUKE [139] Link prediction CN15K DistMult Fact classification NL27K 2021 PPI5K BEURRE [21] Gumbel boxes Confidence prediction CN15K 2021 Relation fact ranking NELL27k PASSLEAF [22] PPI5K RotatE Tail Entity Prediction NL27K ComplEx Confidence Prediction CN15K 2021 DistMult WN18RR FB15K237 GMUC [159] Minimum Similarity Link Prediction NL27K 2021 Confidence Prediction ConfE [163] RESCAL Entity Type Noise Detection FB15kET (synthetic) 2021 Entity Type Prediction YAGO43k (synthetic) FocusE [109] Modifiable Scoring Layer Link Prediction (High-Valued Links) O*NET20K 2021 CN15K NL27K PPI5K WaExt [76] TransE CN15K 2022 TransH Link Prediction NL27K DistMult Triple Classification PPI5K ComplEx Wang et al. [140] Similarity Confidence Prediction CN15K 2022 Tail Entity Prediction NL27K MUKGE [90] Circular correlation Confidence Prediction CN15K 2024 Relation Fact Ranking NL27K Relation Fact Classification PPI5K
@@ 8.2 Fusion Approaches
Fusion methods are identified in Table 3, which provides an overview and indicates if the methods can handle some characteristics of data or data sources such as numerical data, categorical data, data granularity, and dependency between data sources.
Model Task Modeling Awareness Datasets Year TI SQ Numerical Data Categorical Data Granularity Source dependence TruthFinder [157] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ Vagueness $\bullet$ Book authors 2007 ACCU [32] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ Vagueness $\bullet$ Synthetic 2009 LFC [114] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ $\circ$ Digital mammography and Breast MRI 2010 GTM [161] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\circ$ Fuzziness $\circ$ Wikipedia edit history of city population and People biographies 2012 LTM [162] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ Completeness $\circ$ Book author, Movie director, and Synthetic 2012 POPACCU [33] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ $\circ$ Books (from AbeBooks.com) and Flight 2012 LCA [111] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ $\circ$ Books, Population, Stocks, and FantasySCOTUS 2013 CRH [83] $\bullet$ $\bullet$ Optimization problem $\bullet$ $\bullet$ Vagueness $\circ$ Weather Forecast, Stocks, and Flight 2014 CATD [82] $\bullet$ $\bullet$ Optimization $\bullet$ $\bullet$ $\circ$ $\circ$ City Population, Biography, and Indoor Floorplan 2014 KBT [34] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ $\circ$ Triples collected by Knowledge Vault 2015 Synthetic FaitCrowd [91] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ $\circ$ SFV and Dataset from crowdsourcing platform 2015 DOCS[164] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ $\circ$ ItemCompare, 4-Domain, Yahoo QA, and SFV 2016 ASUMS [5] $\bullet$ $\bullet$ Belief functions $\circ$ $\bullet$ Fuzziness $\circ$ Synthetic and People biographies 2016 KDEm [138] $\bullet$ $\bullet$ Optimization $\bullet$ $\circ$ Completeness $\circ$ Synthetic and Population 2016 SLiMFAST [115] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ $\circ$ $\bullet$ Stocks, Demonstrations, Crowds, and Genomics 2017 MDC [87] $\bullet$ $\bullet$ Representation Learning $\circ$ $\bullet$ $\circ$ $\circ$ Dataset created from baobaozhido (crowdsourcing platform) 2017 Optimization problem HYBRID [81] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ Completeness $\circ$ Book and Synthetic 2017 TDH [71] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ Fuzziness $\bullet$ BirthPlaces and Heritages 2019 Record Fusion [55] $\bullet$ $\circ$ Softmax Classifier $\bullet$ $\bullet$ $\circ$ $\circ$ Flight, Stock (1 & 2), Weather, and Address 2020 OKELE [14] $\bullet$ $\bullet$ Probabilistic $\bullet$ $\bullet$ Completeness $\circ$ Synthetic 2020 TKGC [62] $\bullet$ $\bullet$ Representation Learning $\bullet$ $\bullet$ Completeness $\circ$ Dataset built from [14] 2022 Probabilistic Subgraph of Freebase as prior knowledge (KG)
@@
CRH [83] (Conflict Resolution on Heterogeneous Data) estimates the reliability of a source, it uses all types of data simultaneously instead of focusing on a single one. To do this, the authors use a loss function that measures the distance between the unknown truth and the values claimed by the sources for each type of data. To initiate the source reliability score, it first applies a simple conflict resolution method, such as majority or average voting and models the estimation of source and truths weights as an optimization problem.
@@
Then, they use the logarithm to facilitate the computation and to obtain the final scores. Although simple, these definitions integrate the influence of facts where values of different vagueness levels (e.g., “Joe Biden”, “J. Biden”, and “Biden”) can increase the confidence of a fact and take into account the dependence between data sources by a dampening factor acting on the computation of the confidence score.
@@
OKELE [14] models the probability of a fact being true by a latent random variable following a Beta$(\beta_{0},\beta_{1})$ distribution, with $\beta_{1}$ corresponding to prior true count of the fact and $\beta_{0}$ corresponding to its prior false count. The quality of a data source is represented by its error variance $\omega_{s}$ that follows a scaled inverse chi-squared distribution Scale-inv-$\chi^{2}(v_{s},\tau_{s}^{2})$ representing the number of facts provided by the source $v_{s}$ with variance $\tau_{s}^{2}$. The authors argue that this distribution handles the effect of dataset size in the case of long-tail entities. The truth inference is performed by leveraging prior knowledge from existing KGs to identify whether an attribute expects a single or multiple values.